# Bidirectional Information Flow (BIF) - A Sample Efficient Hierarchical Gaussian Process for Bayesian Optimization

### v1

@@ Abstract
Hierarchical Gaussian Process (H-GP) models divide problems into different subtasks, allowing for different models to address each part, making them well-suited for problems with inherent hierarchical structure. However, typical H-GP models do not fully take advantage of this structure, only sending information up or down the hierarchy. This one-way coupling limits sample efficiency and slows convergence. We propose Bidirectional Information Flow (BIF), an efficient H-GP framework that establishes bidirectional information exchange between parent and child models in H-GPs for online training. BIF retains the modular structure of hierarchical models — the parent combines subtask knowledge from children GPs — while introducing top-down feedback to continually refine children models during online learning. This mutual exchange improves sample efficiency, enables robust training, and allows modular reuse of learned subtask models. BIF outperforms conventional H-GP Bayesian Optimization methods, achieving up to $85\%$ and 5x higher $R^{2}$ scores for the parent and children respectively, on synthetic and real-world neurostimulation optimization tasks.
@@ 3.5 Training Bidirectional Information Flow
3:     Generate prior map $p_{s}=\mu_{s}+\gamma\frac{\sigma_{s}}{\sqrt{n_{s}}}\forall s\in S$triangleright S is the set of children. Section 3.2
4:     Create average prior $p_{avg}=\frac{1}{|S|}\sum_{s}^{S}p_{s}$triangleright Refer to Section 3.2
5:     Pass $p_{avg}$ as prior to the mean for Parent $h$triangleright Refer to Section 3.2
6:     Generate Acquisition Map for Parent $a_{h}=\mu_{h}+\kappa\frac{\sigma_{h}}{\sqrt{n_{h}}}$triangleright Refer to Section 3.1
@@
8:     Query environment to get $y_{h}(x_{q})$ as the true environment response
9:     Calculate contributions $c_{s}(x_{q})=\mu_{s}(x_{q})+\kappa\frac{\sigma_{s}(x_{q})}{\sqrt{n_{s}(x_{q})}}\forall s\in S$triangleright Refer to Section 3.3
10:     Partition parent response to children $y_{s}(x_{q})=y_{h}(x_{q})*\frac{e^{c_{s}(x_{q})}}{\sum_{j}^{S}e^{c_{j}(x_{q})}}$
11:     Add respective $x_{q},T[y(x_{q})]$ to each GP’s training set $\triangleright$ Refer to Section 3.4
@@ 4.1 Implementation Details
As can be seen in Algorithm 1, there are a few hyperparameters that must be set for the model to run such as $\kappa,\gamma,r,\text{and, }t$ as well as a few additional hyperparameters inherent to GPBO. We run an independent hyperparameter study to determine the optimal hyperparameters for each application and model. These experiments are presented in the Appendix C. Analyzing the hyperparameter studies, we determine that, in general, the optimal hyperparameters for the BIF pipeline are as follows: $\kappa=9.5,\gamma=3,r=6,\text{and }t=10$ for synthetic datasets and $\kappa=4.0,\gamma=3,r=6,\text{and }t=10$ for the neurostimulation. For the child GPs, we use a Matern Kernel with a lengthscale parameter set to $\nu=0.5$ while the parent uses an additive kernel where the kernels for each children are summed $kernel_{p}=\sum_{s}^{S}kernel_{s}$. For the parameter updates, the ADAM optimizer is used with a learning rate of $\alpha=0.01$ for both the parent and child GP models. The models were run without the use of GPU acceleration and implemented using the GPyTorch Libary (Gardner et al., 2021). Approximate runtime is 1.5 query/s with a maximum budget of 100 queries on Intel 12th Gen i7 processor with 32 Gb of memory.
@@ 4.2 Datasets
Evaluation of the BIF model is done on optimization of three key datasets, each highlighting a different aspect of the model’s performance. The first dataset is a two-dimensional synthetic dataset, where data is generated via a hierarchical process itself: from two children in the hierarchy and one parent. Data is generated through arbitrary functions with a combination in the parent as follows: $f(x)=\frac{sin(2\pi x)}{x-1}+2,g(y)=\frac{sin(2\pi x)}{x-2}+2,\text{and parentfunction }h(x,y)=\frac{f(x)+g(y)}{(x-y)^{2}+\epsilon}$. In this case, the two children GPs aim to optimize functions $f,g$ and feed their knowledge as a prior to the parent. The parent integrates the priors to boost its performance in optimization function $h$. The next dataset used in this study is a higher dimensional 3D dataset with three children and one parent. The functions constituting this optimization set are as follows: $f(x)=-(x-2)^{5},g(y)=sin^{3}(y),k(z)=\frac{log(z+1)+1}{z+1}\text{ and parent function }h(x,y,z)=(f(x)+g(y)+k(z))^{2}+2$. The goal of this dataset is to show the robustness of the BIF architecture to learn higher dimensional tasks. Both synthetic datasets have 10% noise added as described in Section 4.5. The final dataset is the optimization of a neurostimulation dataset with two child targets and one parent generated by combining the stimulation of the two children (Laferriere, 2019). Inputs to the children are the electrode channel on an electrode array implanted in the motor cortex of a non-human primate with the target being maximal muscle activation. This dataset aims to show the applicability of BIF to real-world engineering problems, where success could dictate the implementation of BIF to in-vivo neurostimulation protocols.
@@ 5 Results
Dataset Model Relative Optimum Parent $R^{2}$ Child $R^{2}$ Global AUC
Synthetic 2D Vanilla GPBO 83.51 86.37 - 13,383.17
Laferrière 98.81 92.10 60.70 21,271.38
BIF 96.40 88.26 87.66 22,611.22
Synthetic 3D Vanilla GPBO 44.70 41.36 - 5,068.68
Laferrière 37.67 46.78 8.75 6,000.55
BIF 68.27 86.34 50.74 13,740.73
Neural Stimulation Vanilla GPBO 90.84 82.70 - 14,638.65
Laferrière 98.07 93.81 37.32 18,906.40
BIF 95.99 96.35 71.72 21,910.59
Figure 2: Baseline GPBO models vs. BIF on 2D (a–c) and 3D (d–h) synthetic tasks. (a) shows all model performance for the 2D task; (b-c) show BIF child environment reconstructions (blue: predicted mean/uncertainty, red: ground truth). (d) shows all model performance on the 3D task; (e-g) show BIF child environment reconstructions for 3D, analogous to (b-c). See Section 4.2 for details. Parent Visualizations for 2D available in Appendix D.
@@ 5.1 Synthetic Two Dimensional Dataset
As a first set of analyses, we run our BIF model on the 2D dataset outlined in 4.2 compared to a baseline vanilla GPBO and an algorithm based on Laferrière et al. (2020) on the same dataset with results reported in Table 1 and Figure 2 (a-c). Both Laferrière and BIF models require an initialization period for the children as detailed in Section 4.1 and transfer the information from the children to the parent. Since the Global AUC sums over all query points, quicker learning will result in a higher Global AUC which we see in both hierarchical models. Evidently, the information provided by the children is substantial, allowing the parent model to quickly outperform the vanilla GPBO in the relative optimum score, parent $R^{2}$, and Global AUC of Synthetic 2D in Table 1. We can see both Laferrière et al. (2020) and BIF exhibit similar learning trajectories for the parent models, demonstrating that the changing parameters for the child while learning do not significantly affect the parent’s performance. In contrast to the parent, we see the largest margin of difference with respect to performance when looking at the coefficient of correlation for the children in the finely dotted lines of Figure 2(a). Due to the definition of the baseline Laferrière model, the children do not learn past their initialization sets, freezing their $R^{2}$ score. Observing the training dynamics of the BIF model’s children, we see substantial gains in performance following the random initialization set, where the only information being received is through the contribution method outlined in Section 3.3. These gains in performance highlight the contributions of our work, allowing parallel training of the hierarchy with only one information source. A more in depth interpretation of this difference is visualized in Figure 2 (c-d) where the children’s belief space (blue line) is plotted against the ground truth (red line). We can see the model’s child $R^{2}$ score of 87.66 is sufficient for modeling the environment and accurately discovering the maximum location.
@@ 5.2 Synthetic Three Dimensional Dataset
The three dimensional dataset is where we see the largest performance boost with BIF compared to the two baselines seen in Table 1. As a summary measurement, we refer to the Global AUC metric. In this column of Table 1, we demonstrate that BIF is able to improve the performance of both the other models by over 100%. A near doubling in performance is also seen in both the coefficients of correlation for the parent and children, where the parent $R^{2}$ is 85% higher than the Laferrière model and the child $R^{2}$ improving over fivefold. The RO score is where we see the smallest advantage in performance, with vanilla GPBO ranking second for this metric and being outperformed by 53%. The drastic increases in performance here is an indicator of BIF’s ability to handle complex datasets, far superior than previous baseline models. These metrics for BIF are further described in Figure 2, where the training trajectories are shown. Here we demonstrate the efficiency of our model, with all performance metrics achieving nearly their full performance by 50 queries on a task with 10 times more states than the two dimensional task of Section 5.1. We also emphasize the efficacy of the upwards and downwards information flow outlined in Sections 3.2 and 3.3 as the learning of the children and parent begin much lower than in the two dimensional setting but are still able to extract meaningful information from the environment to train the children. These observations verify the claim of the efficiency of our novel method as well as the efficacy of the contribution system in training the children on a mostly inferred training dataset.
Figure 3: Baseline Laferrière et al. (2020) (a) vs. BIF (b–d) on a real-world neurostimulation task, averaged over 30 runs. (a-b) show overall performance metrics, with BIF achieving higher parent and child $R^{2}$ but lower relative optimum. (c-d) show example BIF child belief contours (blue: prediction; red: ground truth). Parent Visualizations available in Appendix D.
@@ 5.3 Neurostimulation Dataset
As introduced in Section 4.2, all models were tested on a real-world environment of neurostimulation optimization (Laferriere, 2019). The results of this testing are presented in row 3 of Table 1 and Figure 3. Both Laferrière et al. (2020) and BIF outperform Vanilla GPBO in all metrics with BIF achieving the highest $R^{2}$ scores for both parent and child. Notably, BIF outperforms Vanilla GPBO in the parent $R^{2}$ by approximately 14 points and nearly doubles the average child $R^{2}$ of the Laferrière model. These results are visualized in Figure 3 (a-b) as the training curves of Laferrière et al. (2020) and BIF respectively. It is clear that the contribution system presented in Section 3.3 is highly effective, quickly improving the average coefficient of correlation for the children. For a more detailed representation of this metric, Figures 3 (c-d) depict an example child GP representation of the state space following the training protocol. Although the average child $R^{2}$ reaches only 71.72%, the visualization of the predicted vs true environment state spaces shows that the models are accurately depicting the responses of the neurostimulation set as was also seen in the child contours of Figure 2. With these results we verify our claim that our Bidirectional Information Flow method is able to accurately reconstruct the environment of real-world environments with a training dataset inferred by the contribution system approximation. Less noticeable than the child performance is the improvement in the parent’s $R^{2}$ score, improving the performance of the Laferrière model by more than 3% and ultimately a lesser drop in the RO score of the BIF model. The discrepancy in these scores could be attributed to the influence of uncertainty in the RO score, while the $R^{2}$ only uses the mean predictions of the model. We focus on improving the child $R^{2}$ over the RO score due to the novelty of the downward information flow. We prove this tradeoff in scores in Appendix C through the hyperparameter testing of our model showing higher RO scores for lower kappa values and higher child $R^{2}$ with higher kappa.

### v2

@@ Abstract
Hierarchical Gaussian Process (H-GP) models divide problems into different subtasks, allowing for different models to address each part, making them well-suited for problems with inherent hierarchical structure. However, typical H-GP models do not fully take advantage of this structure, only sending information up or down the hierarchy. This one-way coupling limits sample efficiency and slows convergence. We propose Bidirectional Information Flow (BIF), an efficient H-GP framework that establishes bidirectional information exchange between parent and child models in H-GPs for online training. BIF retains the modular structure of hierarchical models — the parent combines subtask knowledge from children GPs — while introducing top-down feedback to continually refine children models during online learning. This mutual exchange improves sample efficiency, enables robust training, and allows modular reuse of learned subtask models. BIF outperforms conventional H-GP Bayesian Optimization methods, achieving up to 4x and 3x higher $R^{2}$ scores for the parent and children respectively, on synthetic and real-world neurostimulation optimization tasks.
@@ 3.5 Training Bidirectional Information Flow
3:  Generate prior map $p_{s}=\mu_{s}+\gamma\frac{\sigma_{s}}{\sqrt{n_{s}}}\forall s\in S$ $\triangleright$ S is the set of children. Section 3.2
4:  Create average prior $p_{avg}=\frac{1}{|S|}\sum_{s}^{S}p_{s}$ $\triangleright$ Refer to Section 3.2
5:  Pass $p_{avg}$ as prior to the mean for Parent $h$ $\triangleright$ Refer to Section 3.2
6:  Generate Acquisition Map for Parent $a_{h}=\mu_{h}+\kappa\frac{\sigma_{h}}{\sqrt{n_{h}}}$ $\triangleright$ Refer to Section 3.1
@@
8:  Query environment to get $y_{p}(x_{q})$ as the true environment response
9:  Calculate contributions $c_{s}(x_{q})=\mu_{s}(x_{q})+\kappa\frac{\sigma_{s}(x_{q})}{\sqrt{n_{s}(x_{q})}}\forall s\in S$ $\triangleright$ Refer to Section 3.3
10:  Partition parent response to children $y_{s}(x_{q})=y_{p}(x_{q})*\frac{e^{c_{s}(x_{q})}}{\sum_{j}^{S}e^{c_{j}(x_{q})}}$
11:  Add $x_{q}\text{ to parent and},T[y(x_{q})]$ to respective child GP’s training set $\triangleright$ Refer to Section 3.4
@@ 4.1 Implementation Details
As can be seen in Algorithm 1, there are a few hyperparameters that must be set for the model to run such as $\kappa,\gamma,r,\text{and, }t$ as well as a few additional hyperparameters inherent to GPBO. We run an independent hyperparameter study to determine the optimal hyperparameters for each application and model. These experiments are presented in the Appendix C. Analyzing the hyperparameter studies, we determine that, in general, the optimal hyperparameters for the BIF pipeline are as follows: $\kappa=7.5,\gamma=3,r=6,\text{and }t=10$ for synthetic datasets and $\kappa=4.0,\gamma=3,r=6,\text{and }t=10$ for the neurostimulation. For the child GPs, we use a Matern Kernel with a lengthscale parameter set to $\nu=0.5$ while the parent uses an additive kernel where the kernels for each children are summed $kernel_{p}=\sum_{s}^{S}kernel_{s}$. For the parameter updates, the ADAM optimizer is used with a learning rate of $\alpha=0.01$ for both the parent and child GP models. The models were run without the use of GPU acceleration and implemented using the GPyTorch Libary (Gardner et al., 2021). Approximate runtime is 1.5 query/s with a maximum budget of 100 queries on Intel 12th Gen i7 processor with 32 Gb of memory. Each experiment was repeated 30 times with randomized initialization sets.
@@ 4.2 Datasets
Evaluation of the BIF model is done on optimization of three key datasets, each highlighting a different aspect of the model’s performance. The first dataset is a two-dimensional synthetic dataset, where data is generated via a hierarchical process itself: from two children in the hierarchy and one parent. Data is generated through arbitrary functions with a combination in the parent as follows: $f(x)=\frac{sin(2\pi x)}{x-1}+2,g(y)=\frac{sin(2\pi y)}{y-2}+2,\text{and parent function }h(x,y)=\frac{f(x)+g(y)}{(x-y)^{2}+\epsilon}$. In this case, the two children GPs aim to optimize functions $f,g$ and feed their knowledge as a prior to the parent. The parent integrates the priors to boost its performance in optimization function $h$. The next dataset used in this study is a higher dimensional 3D dataset with three children and one parent. The functions constituting this optimization set are as follows: $f(x)=-(x-2)^{5},g(y)=sin^{3}(y),k(z)=\frac{log(z+1)+1}{z+1}\text{ and parent function }h(x,y,z)=(f(x)+g(y)+k(z))^{2}+2$. The goal of this dataset is to show the robustness of the BIF architecture to learn higher dimensional tasks. Both synthetic datasets have 10% noise added as described in Section 4.5. The final dataset is the optimization of a neurostimulation dataset with two child targets and one parent generated by combining the stimulation of the two children (Laferriere, 2019). Inputs to the children are the electrode channel on an electrode array implanted in the motor cortex of a non-human primate with the target being maximal muscle activation. This dataset aims to show the applicability of BIF to real-world engineering problems, where success could dictate the implementation of BIF to in-vivo neurostimulation protocols.
@@ 4.3 Child Process Modularity
Dataset Model Relative Optimum Parent $R^{2}$ Child $R^{2}$ Global AUC
Synthetic 2D Vanilla GPBO 89.47 73.25 - 13,462.81
Laferrière 93.02 69.93 54.51 17,389.45
BIF 93.02 70.48 69.26 17,410.21
Synthetic 3D Vanilla GPBO 54.28 13.80 - 4,841.83
Laferrière 44.92 12.24 13.10 5,126.94
BIF 78.60 56.94 48.91 12,197.79
Neural
Stimulation
Vanilla GPBO 92.21 24.55 - 10,275.60
Laferrière 100.00 50.63 36.80 15,166.66
BIF 100.00 73.82 58.68 18,220.20
Figure 2: Baseline GPBO models vs. BIF on 2D (a–c) and 3D (d–g) synthetic tasks. (a) shows all model performance for the 2D task; (b-c) show BIF child environment reconstructions (blue: predicted mean/uncertainty, red: ground truth). (d) shows all model performance on the 3D task; (e-g) show BIF child environment reconstructions for 3D, analogous to (b-c). See Section 4.2 for details. Parent Visualizations for 2D available in Appendix D.
@@ 5.1 Synthetic Two Dimensional Dataset
As a first set of analyses, we run our BIF model on the 2D dataset outlined in 4.2 compared to a baseline vanilla GPBO and an algorithm based on Laferrière et al. (2020) on the same dataset with results reported in Table 1 and Figure 2 (a-c). Both Laferrière and BIF models require an initialization period for the children as detailed in Section 4.1 and transfer the information from the children to the parent. Since the Global AUC sums over all query points, quicker learning will result in a higher Global AUC which we see in both hierarchical models. Evidently, the information provided by the children is substantial, allowing the parent model to quickly outperform the vanilla GPBO in the relative optimum score, and Global AUC of Synthetic 2D in Table 1. We can see both Laferrière et al. (2020) and BIF exhibit similar learning trajectories for the parent models, demonstrating that the actively changing parameters for the child while learning do not significantly affect the parent’s performance. In contrast to the parent, we see the largest margin of difference with respect to performance when looking at the coefficient of correlation for the children in the finely dotted lines of Figure 2(a). Due to the definition of the baseline Laferrière model, the children do not learn past their initialization sets, freezing their $R^{2}$ score. Observing the training dynamics of the BIF model’s children, we see substantial gains in performance following the random initialization set, where the only information being received is through the contribution method outlined in Section 3.3. These gains in performance highlight the contributions of our work, allowing parallel training of the hierarchy with only one information source. A more in depth interpretation of this difference is visualized in Figure 2 (c-d) where the children’s belief space (blue line) is plotted against the ground truth (red line). We can see the model’s child $R^{2}$ score of 69.26 is sufficient for modeling the environment and accurately discovering the maximum location.
@@ 5.2 Synthetic Three Dimensional Dataset
The three dimensional dataset is where we see the largest performance boost with BIF compared to the two baselines seen in Table 1. As a summary measurement, we refer to the Global AUC metric. In this column of Table 1, we demonstrate that BIF is able to improve the performance of both the other models by over 100%. A near doubling in performance is also seen in both the coefficients of correlation for the parent and children, where the parent $R^{2}$ is over four times higher than the Laferrière model and the child $R^{2}$ improving over threefold. The RO score is where we see the smallest advantage in performance, with vanilla GPBO ranking second for this metric and being outperformed by 45%. The drastic increases in performance here is an indicator of BIF’s ability to handle complex datasets, far superior than previous baseline models. These metrics for BIF are further described in Figure 2, where the training trajectories are shown. Here we demonstrate the efficiency of our model, with all performance metrics achieving nearly their full performance by 50 queries on a task with 10 times more states than the two dimensional task of Section 5.1. We also emphasize the efficacy of the upwards and downwards information flow outlined in Sections 3.2 and 3.3 as the learning of the children and parent begin much lower than in the two dimensional setting but are still able to extract meaningful information from the environment to train the children. These observations verify the claim of the efficiency of our novel method as well as the efficacy of the contribution system in training the children on a mostly inferred training dataset.
Figure 3: Baseline Laferrière et al. (2020) (a) vs. BIF (b–d) on a real-world neurostimulation task, averaged over 30 runs. (a-b) show overall performance metrics with highlighted regions showing standard error. BIF achieves higher parent and child $R^{2}$ but lower relative optimum. (c-d) show example BIF child belief contours (blue: prediction; red: ground truth). Parent Visualizations available in Appendix D.
@@ 5.3 Neurostimulation Dataset
As introduced in Section 4.2, all models were tested on a real-world environment of neurostimulation optimization (Laferriere, 2019). The results of this testing are presented in row 3 of Table 1 and Figure 3. Both Laferrière et al. (2020) and BIF outperform Vanilla GPBO in all metrics with BIF achieving the highest $R^{2}$ scores for both parent and child. Notably, BIF outperforms Vanilla GPBO in the parent $R^{2}$ by approximately 8 points and nearly doubles the average child $R^{2}$ of the Laferrière model. These results are visualized in Figure 3 (a-b) as the training curves of Laferrière et al. (2020) and BIF respectively. It is clear that the contribution system presented in Section 3.3 is highly effective, quickly improving the average coefficient of correlation for the children. For a more detailed representation of this metric, Figures 3 (c-d) depict an example child GP representation of the state space following the training protocol. Although the average child $R^{2}$ reaches only 58.68, the visualization of the predicted vs true environment state spaces shows that the models are accurately depicting the responses of the neurostimulation set as was also seen in the child contours of Figure 2. With these results we verify our claim that our Bidirectional Information Flow method is able to accurately reconstruct the environment of real-world environments with a training dataset inferred by the contribution system approximation. Equally notable to the child performance is the improvement in the parent’s $R^{2}$ score, improving the performance of the Laferrière model by more than 45% and both hierarchical models achieving a perfect Relative Optimum score. Higher parent $R^{2}$ scores are possible for the BIF model, however, due to the focus of our research being on the improvement of the children, we select our hyperparameters to optimize the children. We prove this tradeoff in scores in Appendix C through the hyperparameter testing of our model showing higher RO scores for lower kappa values and higher child $R^{2}$ with higher kappa.
@@ 6 Conclusion
7 Acknowledgments
This research was supported by the New Frontiers in Research Fund Exploration (NFRFE-2022-00394), the Natural Sciences and Engineering Research Council of Canada (NSERC; RGPIN-2023-04370), and the TransMedTech Institute (Projet-0131-AI-DBS). Guillaume Lajoie acknowledges support from the NSERC Discovery Grant (RGPIN-2018-04821), the Canada Research Chair in Neural Computations and Interfacing, and a Canada-CIFAR AI Chair. We thank Mila – Quebec AI Institute clusters for its support in providing computational resources for model evaluations. We would also like to thank Numa Dancause for his and his group’s work collecting the previously published non-human primate dataset for which our model is designed.